# Neural Acoustic Multipole Splatting for Room Impulse Response Synthesis

### v1

@@ 1 Introduction
Recently, deep neural networks (DNNs) have gained traction for this task. Initially, DNNs estimated RIRs from given source and receiver positions [10, 11] but lacked physical consistency and struggled with high-frequency components. Other approaches introduced physical constraints to enhance model stability and generalization [12, 13, 14]. They integrated physical priors, like the wave equation, into the loss function to guide learning [12]. For better early reflection predictions, some models used periodic activation functions [13, 15], while others applied the dynamic pulling method to enhance performance in noisy settings [14]. These models need only a few RIRs for training [12] and ensure physical fidelity, yet training can be slow due to explicit PDE residual computation and may face gradient issues, affecting high-frequency accuracy [16].
@@ 2.1 NAMS for Room Impulse Response Synthesis
NAMS generates RIRs by employing multipoles that have adjustable directional patterns and emit signals of limited duration. The directional patterns $D_{p}(f)$ for the $p$-th multipole at frequency $f$ can be expressed in terms of the spherical harmonics [20] as
@@
We model the RIR as the superposition of these multipoles, each of which emits a signal $S_{p}(f)$. With a far-field assumption, an RIR in frequency and time domains can be approximated as
@@
Our model predicts each multipole position $\mathbf{x}_{p}$, its emitting signal $s_{p}(t)=\mathcal{F}^{-1}[S_{p}(f)]$, as well as its spherical harmonic coefficients $b_{nm,p}(t)=\mathcal{F}^{-1}[B_{mn,p}(f)]$, for the given receiver position $\mathbf{x}_{r}$ and inverse Fourier transform $\mathcal{F}^{-1}$. The RIR is then synthesized in the frequency domain from Eq. (2) using a set of estimated parameters $\Theta=\{(\mathbf{x}_{p},S_{p}(f),B_{nm,p}(f))\}$. The model is trained to minimize a loss function designed to compare the ground truth and estimated RIRs (detailed loss functions are described in Section 3.2).
@@
The NAMS architecture to estimate the parameter $\Theta$ is depicted in Fig. 1. The architecture consists of signal and directivity branches for estimating multipole position and signal pairs $\{\mathbf{x}_{p},s_{p}(t)\}_{p=1}^{P}$ and directional patterns $\{D_{p}(f)\}$, respectively. In the upper branch, a set of multipole positions $\{\mathbf{x}_{p}\}$ is declared as learnable network parameters, and only this information is utilized to synthesize the multipole signal $s_{p}(t)$. This process is to ensure that the multipole signals are independent of receiver positions. Specifically, the multipole position is encoded by a positional encoder and then fed into the signal estimation layer (MLP) to generate a short-length signal $s_{p}(t)$ in the time domain. The multipole index $p$ is assigned to the batch dimension, and MLP is trained to generate source signals corresponding to their multipole positions. In the lower branch, the model takes the receiver position $\mathbf{x}_{r}$ as input and subtracts it from multipole positions $\mathbf{x}_{p}$ to calculate the relative positions of multipoles required for deriving $\Omega_{p}(\mathbf{x}_{r})$. The relative multipole positions are also encoded by a positional encoder and concatenated with the encoded $\mathbf{x}_{p}$. From the encoded position vectors, the directivity estimation layer (MLP) synthesizes the spherical harmonic coefficients $\{b_{nm,p}(t)=\mathcal{F}^{-1}[B_{nm,p}(f)]\}$ in the time domain.
@@
The coefficients are then used to generate the directional patterns according to Eq. (1). We use real-valued spherical harmonics and constrain the total energy of $D_{p}(f,\mathbf{x}_{r})$ across frequencies to stabilize the training process. In detail, for a vector $\mathbf{D}_{p}=\left[D_{p}(f_{1},\mathbf{x}_{r}),~\cdots,~D_{p}(f_{F},\mathbf{x}_{r})\right]$ defined for discrete frequencies $f_{1},\cdots,f_{F}$, we apply normalization $\tilde{\mathbf{D}}_{p}={\mathbf{D}_{p}}/{\left\lVert\mathbf{D}_{p}\right\rVert}$ and use it in place of $D_{p}(f,\mathbf{x}_{r})$ of Eq. (2). The entire process described above is differentiable, so we can train both $\{\mathbf{x}_{p}\}_{p=1}^{P}$ and the MLPs via backpropagation. Moreover, the physical constraint on the Helmholtz equation is automatically satisfied, because multipoles constituting Eq. (2) are the solution of the Helmholtz equation.
@@ 2.2 Pruning Neural Multipoles
An excessive number of multipoles can lead to overfitting, significantly degrading computational efficiency. Therefore, it is essential to determine the optimal number of multipoles. To achieve this, we densely distribute the multipoles in the space at model initialization and incorporate pruning stages during training. We regularly perform pruning every 20 epochs after the first 100 epochs of training. The principle of pruning is to remove multipoles with low $s_{p}(t)$ energy. To compute the energy of $s_{p}(t)$ for each multipole, we freeze the model and input a dummy receiver position. Since $s_{p}(t)$ depends solely on $\mathbf{x}_{p}$, the choice of dummy receiver position does not affect the result. Also, the total energy of $D_{p}(f)$ is constrained to one, so we can determine the necessity of each multipole solely by the energy of $s_{p}(t)$. If the calculated energy is below 50% of the global median, the corresponding $\mathbf{x}_{p}$ is removed from the model. After each pruning step, the model re-optimizes $\{\mathbf{x}_{p}\}_{p=1}^{P^{\prime}}$ and estimation layers using the reduced number ($P^{\prime}$) of multipoles. This iterative alternating process allows for the development of a compact NAMS model with less possibility of overfitting.
@@ 3.1 Datasets
Our model’s performance was assessed using both real and synthetic datasets. For the real dataset, we utilized the MeshRIR dataset [21]. As for the synthetic dataset, we conducted simulations of two room environments employing the Treble simulator  [^1] https://www.treble.tech, which offers a hybrid simulation approach integrating wave-based and geometrical acoustics. Specifically, we chose two scenes: Apartment 566 and Apartment 716 from the Treble database. The scenes contain furniture such as sofas, carpets, and beds, with absorption and scattering coefficients taken from the predefined values in the Treble database. Apartments 566 and 716 have volumes of 105 m3 and 183 m3, respectively, and follow a Manhattan layout. The reverberation times (T60) of Apartments 566 and 716, averaged for all receiver positions, were 0.48 s and 1.80 s, respectively. We positioned a single omnidirectional source at a fixed location selected at random. RIRs were simulated for 1,000 randomly chosen receiver positions. Fig. 2 depicts the employed room layouts. Across all experiments, we randomly split the RIRs into training and testing sets with a ratio of 9:1. All RIRs were resampled to the 24 kHz sampling rate and trimmed at 0.1 seconds.
@@ 3.2 Implementation Details
During the initial phase, we densely splatted multipoles in space. The starting locations were arranged as points on a collection of spheres, each centered on the source, with radii incrementing from 1 m to 34 m at intervals of 1 m. On each sphere, 32 points were uniformly distributed using Fibonacci sampling and randomly rotated together. An additional multipole was located at the source position to render the direct sound, yielding a total of 1,089 initial positions. We applied sinusoidal positional embedding with 10 sine–cosine frequencies to encode source and receiver positions, following [10]. A 3-layer MLP with 512 hidden units per layer was used to predict $s_{p}(t)$, and the same architecture was employed to predict $b_{nm,p}(t)$. The output $s_{p}(t)$ has a duration of 3 ms, and $b_{nm,p}(t)$ includes spherical harmonic coefficients up to the 3rd order, yielding a 16-channel output with a duration of 3 ms.
@@
Phase Amp. Env. T60 C50 EDT Phase Amp. Env. T60 C50 EDT Phase Amp. Env. T60 C50 EDT
NAF[10] 1.62 0.57 1.98 3.5 0.88 31.0 1.62 0.77 5.05 15.2 8.66 22.0 1.62 0.76 6.56 21.4 7.35 18.8
AVR[17] 1.05 0.28 1.44 2.9 0.66 19.4 1.62 0.46 4.20 5.0 1.20 29.8 1.62 0.54 5.79 9.0 2.46 24.2
NAMS 0.80 0.11 1.21 2.0 0.34 9.8 1.60 0.22 4.46 3.4 0.48 12.0 1.60 0.30 6.70 6.3 0.82 13.7
Table 1: RIR estimation performance comparison with existing models. The evaluation metrics include the phase error (Phase.), amplitude error (Amp.), envelope error (Env., %), relative T60 error (%), C50 error (dB), and EDT error (ms).
@@
Phase Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$ Phase Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$ Phase Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$
mono. sparse 0.93 0.19 1.36 2.3 0.39 12.4 273 1.7 1.61 0.26 4.81 4.0 0.60 16.0 307 1.7 1.62 0.59 10.25 7.8 0.82 16.1 273 1.7
multi. sparse 0.84 0.15 1.30 2.1 0.37 11.5 273 2.1 1.60 0.22 4.50 3.6 0.43 11.5 307 2.2 1.61 0.40 8.20 5.9 0.76 13.5 273 2.1
mono. dense 0.83 0.15 1.27 2.2 0.37 10.8 1089 2.0 1.61 0.28 4.89 4.0 0.69 17.0 1089 2.0 1.62 0.50 9.33 7.2 0.99 16.4 1089 2.0
multi. dense 0.78 0.12 1.20 2.0 0.33 10.2 1089 4.5 1.60 0.22 4.39 3.9 0.50 13.3 1089 4.6 1.61 0.42 8.47 6.3 0.87 13.6 1089 4.6
multi. dense w/ pruning 0.80 0.11 1.21 2.0 0.34 9.8 225 2.2 1.60 0.22 4.46 3.4 0.48 12.0 276 2.2 1.60 0.30 6.70 6.3 0.82 13.7 240 2.1
Table 2: Efficacy of multipole and pruning. Performance comparison of monopole and multipole splatting. ’mono.’ and ’multi.’ represent monopole and multipole settings, respectively, whereas ’sparse’ and ’dense’ denote the small and large number of poles at initialization.
@@
Table 3: Space and time complexity Comparison of the model’s parameter size (Param.), number of sample points or multipoles (#pts), and inference time ($\mathrm{T}_{\mathrm{Inf}}$, ms).
Fig. 3: Comparison of spatial magnitude distribution. Overall magnitude distribution averaged over the 1/3 octave band centered at 4 KHz (Scene from MeshRIR).
@@
Fig. 5: Multipole positions and signal energies with and without pruning. Circles indicate the multipoles inside the room, with deeper red indicating higher energy. The square denotes the source. (a) dense w/o pruning, (b) sparse w/o pruning, (c) dense w/ pruning.
@@ 4.1 Performance Comparison
We compared our model with NAF [10] and AVR [17] that estimate RIRs only from the given source and receiver positions. For a fair comparison, all models were trained for the same source position in the real and synthetic datasets. The point-neuron model [16] was excluded from comparison because its code is not publicly available. Instead, in our ablation study, we evaluated performance differences between monopole and multipole configurations.
The model performances were assessed using the following metrics: phase error, amplitude error, envelope error, reverberation time (T60) error, clarity (C50) error, and early decay time (EDT) error [17]. Phase and amplitude errors are relative errors with no unit; envelope error is presented as a relative percentage error, and the units for T60, C50, and EDT errors are %, dB, and milliseconds, respectively.
We present the experimental results in Table 1. On the MeshRIR dataset, NAMS outperforms existing methods on all metrics. Similar trends can be seen on the Apartment 566 and 716 datasets, except for the envelope error. However, all models yield a phase error of approximately 1.62 on the synthetic dataset, probably because the rapid phase increase from large propagation delay is wrapped within $\pm\pi$. To visualize the differences in the estimated sound field, we present the spatial magnitude distribution of sound fields estimated from the MeshRIR dataset. The results show a clear difference of NAMS compared to baseline models with less spatial jitter. In terms of RIR waveforms shown in Fig. 4, NAMS captures the direct sound and early reflection peaks more accurately than AVR and NAF.
In Table 3, we also compare the space and time complexity of models. NAMS also shows a fast inference time ($\mathrm{T}_{\mathrm{Inf}}$) of 2.1–2.2 ms, comparable to NAF (1.9–2.0 ms). NAMS achieves this speed using only a few hundred multipoles, whereas AVR requires sampling 204,800 points. In addition, our model has an advantage in terms of parameter size. AVR [17] has 57.2 million parameters, whereas NAMS and NAF [10] have 1.8 million and 2.7 million parameters, respectively. However, the large parameter count of AVR [17] is mainly due to its use of hash grid encoding [24]. Excluding the hash grid encoding, AVR has a slightly higher number of parameters.

### v2

@@ 1 Introduction
Recently, deep neural networks (DNNs) have gained traction for this task. Initially, DNNs estimated RIRs from given source and receiver positions [14, 23] but lacked physical consistency and struggled with high-frequency components. Other approaches introduced physical constraints to enhance model stability and generalization [9, 15, 18, 12]. They integrated physical priors, like the wave equation, into the loss function to guide learning [9, 15]. For better early reflection predictions, some models used periodic activation functions [18, 21], while others applied the dynamic pulling method to enhance performance in noisy settings [12]. These models need only a few RIRs for training [9] and ensure physical fidelity, yet training can be slow due to PDE residual computation, may face gradient issues, and exhibit limited high-frequency accuracy [3].
@@ 2.1 NAMS for Room Impulse Response Synthesis
NAMS generates RIRs by employing multipoles that have adjustable directional patterns and emit signals of limited duration. The directional patterns $D_{p}(f,\mathbf{x}_{r})$ for the $p$-th multipole at frequency $f$ can be expressed in terms of the spherical harmonics [5] as
@@
We model the RIR as the superposition of these multipoles, each of which emits a signal $S_{p}(f)$. With a far-field assumption, the RIR in frequency domain can be approximated as
@@
Our model predicts each multipole position $\mathbf{x}_{p}$, its emitting signal $s_{p}(t)=\mathcal{F}^{-1}[S_{p}(f)]$, as well as its spherical harmonic coefficients $b_{nm,p}(t)=\mathcal{F}^{-1}[B_{nm,p}(f)]$, for the given receiver position $\mathbf{x}_{r}$ and inverse Fourier transform $\mathcal{F}^{-1}$. The RIR is then synthesized in the frequency domain from Eq. (2) using a set of estimated parameters $\Theta=\{(\mathbf{x}_{p},S_{p}(f),B_{nm,p}(f))\}_{p=1}^{P}$. The model is trained to minimize a loss function comparing the ground-truth and estimated RIRs (detailed loss functions are described in Section 3.2).
@@
The NAMS architecture to estimate the parameter set $\Theta$ is depicted in Fig. 1. The architecture consists of signal and directivity branches for estimating multipole position and signal pairs $\{\mathbf{x}_{p},s_{p}(t)\}_{p=1}^{P}$ and directional patterns $\{D_{p}(f,\mathbf{x}_{r})\}_{p=1}^{P}$, respectively. In the upper branch, a set of multipole positions $\{\mathbf{x}_{p}\}$ is declared as learnable network parameters, and only this information is utilized to synthesize the multipole signal $s_{p}(t)$. This process is to ensure that the multipole signals are independent of receiver positions. Specifically, the multipole position is encoded by a positional encoder and then fed into the signal estimation layer (MLP) to generate a short-length signal $s_{p}(t)$ in the time domain. The multipole index $p$ is assigned to the batch dimension, and the MLP is trained to generate source signals corresponding to their multipole positions. In the lower branch, the model takes the receiver position $\mathbf{x}_{r}$ as input and subtracts it from multipole positions $\mathbf{x}_{p}$ to calculate the relative positions of multipoles required for deriving $\mathbf{\Omega}_{p}(\mathbf{x}_{r})$. The relative multipole positions are also encoded by a positional encoder and concatenated with the encoded $\mathbf{x}_{p}$. From the encoded position vectors, the directivity estimation layer (MLP) synthesizes the spherical harmonic coefficients $\{b_{nm,p}(t)=\mathcal{F}^{-1}[B_{nm,p}(f)]\}$ in the time domain.
@@
The coefficients are then used to generate the directional patterns according to Eq. (1). We use real-valued spherical harmonics and constrain the total energy of $D_{p}(f,\mathbf{x}_{r})$ across frequencies to stabilize the training process. In detail, for a vector $\mathbf{D}_{p}=\left[D_{p}(f_{1},\mathbf{x}_{r}),~\cdots,~D_{p}(f_{F},\mathbf{x}_{r})\right]$ defined for discrete frequencies $f_{1},\cdots,f_{F}$, we apply normalization $\tilde{\mathbf{D}}_{p}={\mathbf{D}_{p}}/{\left\lVert\mathbf{D}_{p}\right\rVert}$ and use it in place of $D_{p}(f,\mathbf{x}_{r})$ in Eq. (2). The entire process described above is differentiable, so we can train both $\{\mathbf{x}_{p}\}_{p=1}^{P}$ and the MLPs via backpropagation. Moreover, the physical constraint on the Helmholtz equation is automatically satisfied, because multipoles constituting Eq. (2) are the solution of the Helmholtz equation.
@@ 2.2 Pruning Neural Multipoles
An excessive number of multipoles can lead to overfitting, significantly degrading computational efficiency. Therefore, it is essential to determine the optimal number of multipoles. To achieve this, we densely distribute the multipoles in the space at model initialization and incorporate pruning stages during training. We regularly perform pruning every 20 epochs after the first 100 epochs of training. The principle of pruning is to remove multipoles with low $s_{p}(t)$ energy. To compute the energy of $s_{p}(t)$ for each multipole, we freeze the model and input a dummy receiver position. Since $s_{p}(t)$ depends solely on $\mathbf{x}_{p}$, the choice of dummy receiver position does not affect the result. Also, the total energy of $D_{p}(f,\mathbf{x}_{r})$ is constrained to one, so we can determine the necessity of each multipole solely by the energy of $s_{p}(t)$. If the calculated energy is below 50% of the global median, the corresponding $\mathbf{x}_{p}$ is removed from the model. After each pruning step, the model re-optimizes $\{\mathbf{x}_{p}\}_{p=1}^{P^{\prime}}$ and estimation layers using the reduced number ($P^{\prime}$) of multipoles. This iterative alternating process allows for the development of a compact NAMS model with less possibility of overfitting.
@@ 3.1 Datasets
Our model’s performance was assessed using both real and synthetic datasets. For the real dataset, we utilized the MeshRIR dataset [11]. As for the synthetic dataset, we conducted simulations of two room environments employing the Treble simulator  [^1] https://www.treble.tech, which offers a hybrid simulation approach integrating wave-based and geometrical acoustics. Specifically, we chose two scenes: Apartment 566 and Apartment 716 from the Treble database. The scenes contain furniture such as sofas, carpets, and beds, with absorption and scattering coefficients taken from the predefined values in the Treble database. Apartments 566 and 716 have volumes of 105 m3 and 183 m3, respectively, and follow a Manhattan layout. The reverberation times (T60) of Apartments 566 and 716, averaged over all receiver positions, were 0.48 s and 1.80 s, respectively. We positioned a single omnidirectional source at a fixed location selected at random. RIRs were simulated for 1,000 randomly chosen receiver positions. Fig. 2 depicts the employed room layouts. Across all experiments, we randomly split the RIRs into training and testing sets with a ratio of 9:1. All RIRs were resampled to 24 kHz sampling rate and trimmed to 0.1 seconds.
@@ 3.2 Implementation Details
During the initial phase, we densely splatted multipoles in space. The starting locations were arranged as points on a collection of spheres, each centered on the source, with radii incrementing from 1 m to 34 m at intervals of 1 m. On each sphere, 32 points were uniformly distributed using Fibonacci sampling and randomly rotated together. An additional multipole was located at the source position to render the direct sound, yielding a total of 1,089 initial positions. We applied sinusoidal positional embeddings with 10 sine–cosine frequencies to encode source and receiver positions, following [14]. A 3-layer MLP with 512 hidden units per layer was used to predict $s_{p}(t)$, and the same architecture was employed to predict $b_{nm,p}(t)$. The output $s_{p}(t)$ has a duration of 3 ms, and $b_{nm,p}(t)$ includes spherical harmonic coefficients up to the 3rd order, yielding a 16-channel output with a duration of 3 ms.
@@
Amp. Env. T60 C50 EDT Amp. Env. T60 C50 EDT Amp. Env. T60 C50 EDT
NAF[14] 0.57 1.98 3.5 0.88 31.0 0.77 5.05 15.2 8.66 22.0 0.76 6.56 21.4 7.35 18.8
AVR[13] 0.28 1.44 2.9 0.66 19.4 0.46 4.20 5.0 1.20 29.8 0.54 5.79 9.0 2.46 24.2
NAMS 0.11 1.21 2.0 0.34 9.8 0.22 4.46 3.4 0.48 12.0 0.30 6.70 6.3 0.82 13.7
Table 1: RIR estimation performance comparison with existing models. The evaluation metrics include the amplitude error (Amp.), envelope error (Env., %), relative T60 error (%), C50 error (dB), and EDT error (ms).
@@
Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$ Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$ Amp. Env. T60 C50 EDT #pts $\mathrm{T}_{\mathrm{Inf}}$
mono. sparse 0.19 1.36 2.3 0.39 12.4 273 1.7 0.26 4.81 4.0 0.60 16.0 307 1.7 0.59 10.25 7.8 0.82 16.1 273 1.7
multi. sparse 0.15 1.30 2.1 0.37 11.5 273 2.1 0.22 4.50 3.6 0.43 11.5 307 2.2 0.40 8.20 5.9 0.76 13.5 273 2.1
mono. dense 0.15 1.27 2.2 0.37 10.8 1089 2.0 0.28 4.89 4.0 0.69 17.0 1089 2.0 0.50 9.33 7.2 0.99 16.4 1089 2.0
multi. dense 0.12 1.20 2.0 0.33 10.2 1089 4.5 0.22 4.39 3.9 0.50 13.3 1089 4.6 0.42 8.47 6.3 0.87 13.6 1089 4.6
multi. dense w/ pruning 0.11 1.21 2.0 0.34 9.8 225 2.2 0.22 4.46 3.4 0.48 12.0 276 2.2 0.30 6.70 6.3 0.82 13.7 240 2.1
Table 2: Effectiveness of multipole and pruning. Performance comparison of monopole and multipole splatting. ’mono.’ and ’multi.’ represent monopole and multipole settings, respectively, whereas ’sparse’ and ’dense’ denote the small and large number of poles at initialization.
@@
Table 3: Space and time complexity. Comparison of the model parameter size (Param.), number of sample points or multipoles (#pts), and inference time ($\mathrm{T}_{\mathrm{Inf}}$, ms).
Fig. 3: Comparison of spatial magnitude distributions. Overall magnitude distribution averaged over the 1/3 octave band centered at 4 kHz (Scene from MeshRIR).
@@
(a) Dense w/o pruning
(b) Sparse w/o pruning
(c) Dense w/ pruning
Fig. 5: Multipole positions and signal energies with and without pruning. Circles indicate the multipoles inside the room, with darker red indicating higher energy. The black square denotes the source, and the gray dashed square indicates the measurement region.
@@ 4.1 Performance Comparison
We compared our model with NAF [14] and AVR [13] that estimate RIRs only from the given source and receiver positions. For a fair comparison, all models were trained for the same source position in the real and synthetic datasets. The point-neuron model [3] was excluded from the comparison because its code is not publicly available. Instead, in our ablation study, we evaluated performance differences between monopole and multipole configurations.
The model performances were assessed using the following metrics: amplitude error, envelope error, reverberation time (T60) error, clarity (C50) error, and early decay time (EDT) error [13]. Amplitude error is relative error with no unit; envelope error is presented as a relative percentage error, and the units for T60, C50, and EDT errors are %, dB, and milliseconds, respectively.
We present the experimental results in Table 1. On the MeshRIR dataset, NAMS outperforms existing methods on all metrics. Similar trends can be seen on the Apartment 566 and 716 datasets, except for the envelope error. To visualize the differences in the estimated sound field, we present the spatial magnitude distribution of sound fields estimated from the MeshRIR dataset. The results show a clear difference of NAMS compared to baseline models with less spatial jitter. In terms of RIR waveforms shown in Fig. 4, NAMS captures the direct sound and early reflection peaks more accurately than AVR and NAF.
In Table 3, we also compare the space and time complexity of models. NAMS also shows a fast inference time ($\mathrm{T}_{\mathrm{Inf}}$) of 2.1–2.2 ms, comparable to NAF (1.9–2.0 ms). NAMS achieves this speed using only a few hundred multipoles, whereas AVR requires sampling 204,800 points. In addition, our model has an advantage in terms of parameter size. AVR [13] has 57.2 million parameters, whereas NAMS and NAF [14] have 1.8 million and 2.7 million parameters, respectively. However, the large parameter count of AVR [13] is mainly due to its use of hash grid encoding [17]. Excluding the hash grid encoding, AVR has 2.0 million parameters, slightly higher than NAMS.